# SiamGPT: Quality-First Fine-Tuning for Stable Thai Text Generation

### v1

@@ Abstract
Open-weights large language models remain difficult to deploy for Thai due to unstable generation under complex instructions, despite strong English performance. To mitigate these limitations, We present SiamGPT-32B, an open-weights model based on Qwen3-32B, fine-tuned with a Quality-First strategy emphasizing curated supervision over data scale. The fine-tuning pipeline combines translated high-complexity English instruction data with a Thai-adapted AutoIF framework for instruction and linguistic constraints. Using supervised fine-tuning only, without continual pretraining or corpus expansion, SiamGPT-32B improves instruction adherence, multi-turn robustness, and linguistic stability. Evaluations on the SEA-HELM benchmark show that SiamGPT-32B achieves the strongest overall performance among similar-scale open-weights Thai models, with consistent gains in instruction following, multi-turn dialogue, and natural language understanding.
@@ 1 Introduction
Although open-weights large language models (LLMs) such as Gemma, Qwen, Deepseek, and Mistral (gemmateam2025gemma3technicalreport; yang2025qwen3technicalreport; deepseekai2025deepseekv3technicalreport; mistralai2025magistral) achieve strong performance on English benchmarks without additional fine-tuning, their Thai-language performance, while non-trivial, remains noticeably weaker and more sensitive to instruction complexity without further adaptation. Consequently, recent Thai-centric efforts such as Typhoon (pipatanakul2023typhoonthailargelanguage; pipatanakul2024typhoon2) and OpenThaiGPT (OTG) (yuenyong2024openthaigpt15; yuenyong2025openthaigpt16r1thaicentric) focus on adapting multilingual base models through various fine-tuning approaches to improve Thai benchmark performance.
While these efforts substantially improved Thai language understanding and benchmark performance, generation-time issues such as code-switching, instruction sensitivity, and multi-turn inconsistency can still be observed in Thai text generation. These limitations are particularly problematic in production settings where models act as final response generators within agentic or tool-augmented systems and must reliably synthesize upstream outputs into fluent Thai (plaat2025agenticlargelanguagemodels; zhang2025agenticcontextengineeringevolving; zhang2025deepanalyzeagenticlargelanguage).
@@
Our approach follows a Quality-First design philosophy (li2024quantityqualityboostingllm) that prioritizes carefully curated supervision over data scale. Specifically, we translate high-complexity English instruction datasets into Thai using gemma-3-27b-it [^5] https://huggingface.co/google/gemma-3-27b-it, the highest-performing model on the Southeast Asian Holistic Evaluation of Language Models (SEA-HELM) Thai translation benchmark (leong2025seahelm). We further incorporate a Thai-adapted AutoIF framework (dong2024autoif) to enforce instruction and linguistic constraints through deterministic verification. This pipeline produces a compact, high-fidelity training corpus that is used for supervised fine-tuning of SiamGPT-32B, directly targeting generation stability, instruction adherence, and multi-turn consistency.
@@ 2.2 Data Curation Pipeline
To operationalize the Quality-First strategy, we construct a dual-stream data curation pipeline illustrated in Figure 2. The first stream transfers reasoning structure and conversational patterns by translating high-quality English instruction datasets into Thai using a high-fidelity LLM. The second stream enforces strict instruction-following behavior and Thai-specific linguistic constraints through a Thai-adapted AutoIF framework. The outputs of both streams are merged into a compact, high-fidelity corpus used for supervised fine-tuning.
Figure 2: SiamGPT data curation pipeline. The framework follows a dual-stream design: (top) an LLM-based translation pipeline that converts high-quality English instruction datasets into Thai using gemma-3-27b-it; (bottom) a Thai-adapted AutoIF pipeline that generates and verifies instruction-following data from Thai seed instructions and executable test code. Outputs from both streams are merged into the final training corpus for supervised fine-tuning.
2.2.1 Translation-Based Instruction Transfer
High-quality English instruction datasets provide dense supervision for reasoning, dialogue control, and instruction adherence, but comparable Thai resources remain limited. We therefore translate selected open-source English instruction datasets into Thai, ensuring that complex reasoning traces and multi-turn structure are preserved.
To select an appropriate translation model, we leverage SEA-HELM benchmark (susanto2025seahelm), focusing on the Thai Natural Language Generation translation metric. Our analysis shows that Gemma-3-27b-it achieves the highest Thai translation score (93.03). Accordingly, we adopt Gemma-3-27b-it as the translation backbone because of its strong handling of Thai grammar and contextual nuance, as well as its favorable trade-off between translation quality and computational efficiency compared to substantially larger models in the 70B-parameter regime.
@@ 2.2.2 Thai-Adapted AutoIF Constraint Enforcement
While translation transfers reasoning structure, it is insufficient for enforcing strict instruction adherence and Thai orthographic correctness. To address this gap, we adapt the AutoIF framework to the Thai linguistic context.
AutoIF replaces subjective human annotation with deterministic, programmatic verification by validating model outputs against explicit constraints using executable scripts. We retain the original English AutoIF seed instructions to preserve universal instruction-following behavior and mitigate catastrophic forgetting (luo2025empiricalstudycatastrophicforgetting).
In addition, we introduce a custom Thai AutoIF component consisting of 39 manually curated Thai seed instructions. Following prior evidence that carefully designed, language-specific supervision improves Thai dataset quality and controllability (pipatanakul2024typhoon2), these instructions target Thai-specific phenomena such as vowel placement rules, consonant usage constraints, and controlled lexical forms. This dual-constraint design preserves high-level reasoning capabilities while enabling precise control over Thai text generation.
@@ 2.3 Final Corpus Construction and Ablation Findings
We initially explored broad-spectrum training recipes, including large mixed Supervised Fine-Tuning (SFT) corpora, Direct Preference Optimization (DPO), and hybrid SFT+DPO pipelines. Extensive ablation studies revealed that increasing data volume degraded generation stability and coherence, and that DPO integration yielded marginal gains in instruction-following metrics while causing regressions in other performance dimensions (li2025limrrlscaling).
Based on these findings, we adopt a minimalist, Pareto-optimal corpus composed of two high-fidelity sources: QuixiAI/SystemChat-2.0 (translated) and AutoIF (English + Thai). This compact corpus prioritizes stability, instruction adherence, and multi-turn consistency over coverage, and forms the sole supervision used to fine-tune SiamGPT-32B.
@@ 3.1 Training Corpus
SystemChat-2.0 (Translated). [^6] https://huggingface.co/datasets/siamaids/SiamAI_Systemchat2.0_Thai We translated 141,000 multi-turn conversations from the QuixiAI/SystemChat-2.0 dataset(systemchat2024). The dataset contains system-prompted dialogues ranging from 3 to 25 turns per conversation. All translations were performed using gemma-3-27b-it, as described in Section 2.2.1.
@@ 3.2 Training Framework and Infrastructure
We initialize training from the Qwen3-32B model. Fine-tuning is conducted using Volcano Engine Reinforcement Learning (VeRL)(Sheng_2025), a high-performance post-training framework that supports flexible dataflow composition and integrates Fully Sharded Data Parallelism (FSDP) with Flash Attention.
@@ 3.3 Optimization and Hyperparameters
We perform supervised fine-tuning (SFT) using the high-fidelity corpus described above. Following the post-training protocol established for Qwen2.5 (qwen2025qwen25technicalreport), we omit learning-rate warmup, as prior work shows that instruction-tuned checkpoints benefit from immediate exposure to the target learning rate rather than gradual warmup.
@@ 3.4 Evaluation Benchmarks
Instruction Following (SEA-IFEval). Instruction adherence is evaluated using SEA-IFEval, a Thai-localized adaptation of IF-Eval (zhou2023ifeval), translated and culturally adapted by native Thai speakers. The benchmark measures compliance with explicit formatting and content constraints (susanto2025seahelm).
Multi-Turn Dialogue (SEA-MTBench). Multi-turn conversational ability is assessed using Thai-MTBench, developed by VISTEC (payoungkhamdee2024mtbenchthai), following the LLM-as-a-Judge paradigm (zheng2023mtbench). The benchmark consists of 68 Thai-translated prompts and uses GPT-4o (gpt-4o-2024-05-13) as the judge model.
Thai Exam (ThaiExam). Thai-specific knowledge and reasoning are evaluated using ThaiExam (thaiexam2024), a multiple-choice benchmark covering ONET, TGAT, TPAT-1, A-Level, and IC licensing exams.
@@ 4.1 Impact of Quality-First Fine-Tuning on Qwen3-32B
The largest improvement is observed in multi-turn dialogue robustness. Performance on SEA-MTBench increases from 57.94 to 75.81, indicating that translated SystemChat-2.0 supervision substantially improves contextual consistency across turns. This shift reflects a transition from single-turn completion behavior to more reliable multi-turn conversational interaction.
@@
Overall, the average score across benchmarks increases from 68.46 to 76.83, confirming that Quality-First fine-tuning yields consistent improvements across stability, instruction following, dialogue robustness, and language understanding.
@@
NLU (QA + Sentiment) 59.80 67.95 (+8.15)
Average 68.46 76.83 (+8.37)
Table 2: Impact of Quality-First fine-tuning on the Qwen3-32B base model. SiamGPT-32B improves performance across all benchmarks, with the largest gains in multi-turn dialogue and instruction following.
@@ 4.2 SEA-HELM Leaderboard Comparison
We further compare SiamGPT-32B against Thai open-weights models of comparable scale, including Typhoon2.5-Instruct (30B) (pipatanakul2024typhoon2) and OTG-R1 (32B) (yuenyong2025openthaigpt16r1thaicentric). Results are shown in Table 3.
@@
Instruction Following (SEA-IFEval) 83.00 79.00 54.00
Safety (Toxicity Detection) 44.19 29.68 41.42
Multi-Turn Dialogue (SEA-MTBench) 75.81 76.16 59.69
Average 63.59 60.44 55.28
Table 3: SEA-HELM benchmark comparison among Thai open-weights models in the 30B–32B class. Average denotes the mean score across all reported benchmarks.
@@ 5 Discussion
Factuality and grounding: Like other LLMs, SiamGPT-32B can hallucinate or produce factually incorrect claims, especially in open-domain settings or when prompts imply missing context huang2023hallucination. This is particularly relevant because our intended deployment is as a final response synthesizer in agentic workflows: the quality and truthfulness of the final answer are bounded by the quality of upstream retrieval, tool outputs, and intermediate reasoning traces. Retrieval-augmented generation can reduce unsupported claims by grounding responses in external evidence, but it is not a complete fix and still requires careful system design and verification lewis2020rag; huang2023hallucination.
@@
Decoding pathologies (repetition and looping): Neural text generation is known to exhibit degeneration behaviors, such as repetition, loss of coherence, and looping, depending on the decoding strategy and prompt structure holtzman2020degeneration. Consistent with this literature, our internal testing indicates that SiamGPT-32B may loop when used as a standalone creative writer rather than within its intended agentic pipeline. In practice, this suggests deploying SiamGPT with (i) stronger context anchoring, (ii) decoding controls (e.g., repetition penalties / sampling constraints), and (iii) refusal or stop criteria for runaway generations holtzman2020degeneration; li2020unlikelihood.
Translation-induced artifacts and cultural nuance: A core design choice in our Quality-First strategy is translating high-complexity English instruction datasets into Thai. While this transfers reasoning structure, translation-based supervision can introduce “translationese” artifacts (unnatural phrasing or distribution shift versus native text), potentially affecting fluency and stylistic naturalness in downstream generations zhang2019translationese. This may partially relate to the SEA-HELM pattern where SiamGPT-32B achieves strong overall performance but lags behind comparable models on NLG. Future iterations should increase the proportion of native Thai instruction and conversational data and explicitly optimize for Thai naturalness while preserving stability.
@@
Evaluation coverage and metric blind spots: Our stability metric operationalizes instability as emitting non-Thai characters under Thai prompts. While this captures a critical production failure mode, it does not fully represent other types of instability (e.g., subtle romanization, punctuation mixing, dialectal variation, or semantic drift). The metric is already designed to allow legitimate mixed tokens and to explain the remaining blind spots (romanization, punctuation mixing, dialect, long-context drift). In addition, widely used instruction and dialog benchmarks capture only slices of real-world behavior: MT-Bench is informative for multi-turn coherence but is also sensitive to judge/model biases, while instruction-following metrics emphasize constraint compliance rather than factual correctness zheng2023mtbench; zhou2023ifeval. Broader evaluation should include human preference tests with Thai speakers, stress tests over long contexts, and targeted domain benchmarks for intended deployments.
@@
Security and misuse risks: When integrated into tool-using or agentic systems, LLMs are vulnerable to prompt injection and related attacks that attempt to override system intent or exfiltrate sensitive information owasp2025top10; lliu2024promptinjection. For safety-critical deployments, we recommend standard secure-by-design measures (least-privilege tool access, input/output filtering, and monitoring) and adversarial testing. Automated red-teaming benchmarks provide a more systematic way to measure the robustness of refusal and safety behavior mazeika2024harmbench.
@@
Data contamination and evaluation leakage: Because our training corpus is derived from public sources and machine translations of existing instruction datasets, there is a non-zero risk of overlap with downstream evaluation items or closely related paraphrases. While we rely on the official SEA-HELM pipeline for standardized comparison, future work should include systematic decontamination checks (e.g., near-duplicate detection between training data and benchmark prompts) to strengthen the validity of reported gains.

### v3

@@ Abstract
Open-weights large language models remain difficult to deploy for Thai due to unstable generation under complex instructions, despite strong English performance. To mitigate these limitations, We present SiamGPT-32B, an open-weights model based on Qwen3-32B, fine-tuned with a Quality-First strategy emphasizing curated supervision over data scale. The fine-tuning pipeline combines high-complexity English instruction data with a Thai-adapted AutoIF framework for instruction and linguistic constraints. Using supervised fine-tuning only, without continual pretraining or corpus expansion, SiamGPT-32B improves instruction adherence, multi-turn robustness, and linguistic stability. Evaluations on the SEA-HELM benchmark show that SiamGPT-32B achieves the strongest overall performance among similar-scale open-weights Thai models, with consistent gains in instruction following, multi-turn dialogue, and natural language understanding.
@@ 1 Introduction
Although open-weights large language models (LLMs) such as Gemma, Qwen, Deepseek, and Mistral (Gemma, 2025; Qwen, 2025; DeepSeek-AI, 2025; Mistral-AI, 2025) achieve strong performance on English benchmarks without additional fine-tuning, their Thai-language performance, while non-trivial, remains noticeably weaker and more sensitive to instruction complexity without further adaptation. Consequently, recent Thai-centric efforts such as Typhoon (Pipatanakul et al., 2023, 2024) and OpenThaiGPT (OTG) (Yuenyong et al., 2024, 2025) focus on adapting multilingual base models through various fine-tuning approaches to improve Thai benchmark performance.
While these efforts substantially improved Thai language understanding and benchmark performance, generation-time issues such as code-switching, instruction sensitivity, and multi-turn inconsistency can still be observed in Thai text generation. These limitations are particularly problematic in production settings where models act as final response generators within agentic or tool-augmented systems and must reliably synthesize upstream outputs into fluent Thai (Plaat et al., 2025; Zhang et al., 2025a, b).
@@
Our approach follows a Quality-First design philosophy (Li et al., 2024) that prioritizes carefully curated supervision over data scale. Specifically, we fine-tune the model using high-complexity English instruction-following data, focusing on instruction adherence and multi-turn interaction. We employ an AutoIF framework adapted to incorporate Thai-specific linguistic and formatting constraints, which are enforced through deterministic verification during supervised fine-tuning. This pipeline produces a compact, high-fidelity training corpus that is used for supervised fine-tuning of SiamGPT-32B, directly targeting generation stability, instruction adherence, and multi-turn consistency.
@@ 2.2 Data Curation Pipeline
To operationalize the Quality-First strategy, we construct a dual-stream data curation pipeline illustrated in Figure 2. The first stream leverages high-quality english instruction-following datasets to transfer reasoning structures and conversational patterns, while mitigating catastrophic forgetting during fine-tuning (Luo et al., 2025). The second stream enforces strict instruction-following behavior and Thai-specific linguistic and formatting constraints through a Thai-adapted AutoIF framework. The outputs of both streams are merged into a compact, high-fidelity corpus used for supervised fine-tuning.
Figure 2: SiamGPT data curation pipeline. The pipeline follows a dual-stream design: (top) a high-quality English instruction-following data stream that provides diverse reasoning structures and multi-turn conversational patterns for supervised fine-tuning; (bottom) a Thai-adapted AutoIF pipeline that enforces instruction-following behavior and Thai-specific linguistic and formatting constraints through deterministic verification. Outputs from both streams are merged into a compact, high-fidelity training corpus used for supervised fine-tuning.
2.2.1 High-quality English Dataset
High-quality English instruction-following datasets provide dense supervision for reasoning, dialogue control, and instruction adherence. In this work, we leverage such datasets to supply complex reasoning structures and multi-turn conversational patterns during supervised fine-tuning. This supervision helps stabilize instruction-following behavior and mitigates catastrophic forgetting when adapting the model to downstream usage scenarios (Luo et al., 2025).
@@ 2.2.2 Thai-Adapted AutoIF Constraint Enforcement
While high-quality english dataset transfers reasoning structure, it is insufficient for enforcing strict instruction adherence and Thai orthographic correctness. To address this gap, we adapt the AutoIF framework to the Thai linguistic context.
AutoIF replaces subjective human annotation with deterministic, programmatic verification by validating model outputs against explicit constraints using executable scripts. We retain the original English AutoIF seed instructions to preserve universal instruction-following behavior and mitigate catastrophic forgetting (Luo et al., 2025).
In addition, we introduce a custom Thai AutoIF component consisting of 39 manually curated Thai seed instructions. Following prior evidence that carefully designed, language-specific supervision improves Thai dataset quality and controllability (Pipatanakul et al., 2024), these instructions target Thai-specific phenomena such as vowel placement rules, consonant usage constraints, and controlled lexical forms. This dual-constraint design preserves high-level reasoning capabilities while enabling precise control over Thai text generation.
@@ 2.3 Final Corpus Construction and Ablation Findings
We initially explored broad-spectrum training recipes, including large mixed Supervised Fine-Tuning (SFT) corpora, Direct Preference Optimization (DPO), and hybrid SFT+DPO pipelines. Extensive ablation studies revealed that increasing data volume degraded generation stability and coherence, and that DPO integration yielded marginal gains in instruction-following metrics while causing regressions in other performance dimensions (Li et al., 2025).
Based on these findings, we adopt a minimalist, Pareto-optimal corpus composed of two high-fidelity sources: QuixiAI/SystemChat-2.0 and AutoIF (English + Thai). This compact corpus prioritizes stability, instruction adherence, and multi-turn consistency over coverage, and forms the sole supervision used to fine-tune SiamGPT-32B.
@@ 3.1 Training Corpus
SystemChat-2.0.555https://huggingface.co/datasets/QuixiAI/SystemChat-2.0 The dataset contains system-prompted dialogues ranging from 3 to 25 turns per conversation.
@@ 3.2 Training Framework and Infrastructure
We initialize training from the Qwen3-32B model. Fine-tuning is conducted using Volcano Engine Reinforcement Learning (VeRL)(Sheng et al., 2025), a high-performance post-training framework that supports flexible dataflow composition and integrates Fully Sharded Data Parallelism (FSDP) with Flash Attention.
@@ 3.3 Optimization and Hyperparameters
We perform supervised fine-tuning (SFT) using the high-fidelity corpus described above. Following the post-training protocol established for Qwen2.5 (Qwen et al., 2025), we omit learning-rate warmup, as prior work shows that instruction-tuned checkpoints benefit from immediate exposure to the target learning rate rather than gradual warmup.
@@ 3.4 Evaluation Benchmarks
Instruction Following (SEA-IFEval). Instruction adherence is evaluated using SEA-IFEval, a Thai-localized adaptation of IF-Eval (Zhou et al., 2023), translated and culturally adapted by native Thai speakers. The benchmark measures compliance with explicit formatting and content constraints (Susanto et al., 2025).
Multi-Turn Dialogue (SEA-MTBench). Multi-turn conversational ability is assessed using Thai-MTBench, developed by VISTEC (Payoungkhamdee et al., 2024), following the LLM-as-a-Judge paradigm (Zheng et al., 2023). The benchmark consists of 68 Thai-translated prompts and uses GPT-4o (gpt-4o-2024-05-13) as the judge model.
Thai Exam (ThaiExam). Thai-specific knowledge and reasoning are evaluated using ThaiExam (Stanford CRFM, 2024), a multiple-choice benchmark covering ONET, TGAT, TPAT-1, A-Level, and IC licensing exams.
Safety. Model safety is evaluated through a toxicity detection task. Models are assessed on their ability to classify Thai text as clean, abusive, or hateful.
@@ 4.1 Impact of Quality-First Fine-Tuning on Qwen3-32B
The largest improvement is observed in multi-turn dialogue robustness. Performance on SEA-MTBench increases from 57.94 to 75.81, indicating that SystemChat-2.0 supervision substantially improves contextual consistency across turns. This shift reflects a transition from single-turn completion behavior to more reliable multi-turn conversational interaction.
@@
Overall, the average score across benchmarks increases from 68.46 to 76.03, confirming that Quality-First fine-tuning yields consistent improvements across stability, instruction following, dialogue robustness, and language understanding.
@@
Code Switching (Stability) 87.70 90.40 (+2.70)
Average 68.46 76.03 (+7.57)
Table 2: Impact of Quality-First fine-tuning on the Qwen3-32B base model. The table follows the order of the visual comparison: IF-Eval, MT-Bench, NLU, Thai Exam, and Code Switching.
@@ 4.2 SEA-HELM Leaderboard Comparison
We further compare SiamGPT-32B against Thai open-weights models of comparable scale, including Typhoon2.5-Instruct (30B) (Pipatanakul et al., 2024) and OTG-R1 (32B) (Yuenyong et al., 2025). Results are shown in Table 3.
@@
Instruction Following (SEA-IFEval) 83.00 79.00 54.00
Multi-Turn Dialogue (SEA-MTBench) 75.81 76.16 59.69
NLG (Translation + Summarization) 42.06 56.70 54.31
@@
Safety 44.19 29.68 41.42
Average 63.60 60.44 55.78
Table 3: SEA-HELM benchmark comparison among Thai open-weights models in the 30B–32B class. The rows are ordered to match the visual breakdown: IF-Eval, MT-Bench, NLG, NLU, NLR, and Safety.
@@ 5 Discussion
Factuality and grounding: Like other LLMs, SiamGPT-32B can hallucinate or produce factually incorrect claims, especially in open-domain settings or when prompts imply missing context Huang et al. (2023). This is particularly relevant because our intended deployment is as a final response synthesizer in agentic workflows: the quality and truthfulness of the final answer are bounded by the quality of upstream retrieval, tool outputs, and intermediate reasoning traces. Retrieval-augmented generation can reduce unsupported claims by grounding responses in external evidence, but it is not a complete fix and still requires careful system design and verification Lewis et al. (2020); Huang et al. (2023).
@@
Decoding pathologies (repetition and looping): Neural text generation is known to exhibit degeneration behaviors, such as repetition, loss of coherence, and looping, depending on the decoding strategy and prompt structure Holtzman et al. (2019). Consistent with this literature, our internal testing indicates that SiamGPT-32B may loop when used as a standalone creative writer rather than within its intended agentic pipeline. In practice, this suggests deploying SiamGPT with (i) stronger context anchoring, (ii) decoding controls (e.g., repetition penalties / sampling constraints), and (iii) refusal or stop criteria for runaway generations Holtzman et al. (2019); Li et al. (2020).
@@
Evaluation coverage and metric blind spots: Our stability metric operationalizes instability as emitting non-Thai characters under Thai prompts. While this captures a critical production failure mode, it does not fully represent other types of instability (e.g., subtle romanization, punctuation mixing, dialectal variation, or semantic drift). The metric is already designed to allow legitimate mixed tokens and to explain the remaining blind spots (romanization, punctuation mixing, dialect, long-context drift). In addition, widely used instruction and dialog benchmarks capture only slices of real-world behavior: MT-Bench is informative for multi-turn coherence but is also sensitive to judge/model biases, while instruction-following metrics emphasize constraint compliance rather than factual correctness Zheng et al. (2023); Zhou et al. (2023). Broader evaluation should include human preference tests with Thai speakers, stress tests over long contexts, and targeted domain benchmarks for intended deployments.
@@
Security and misuse risks: When integrated into tool-using or agentic systems, LLMs are vulnerable to prompt injection and related attacks that attempt to override system intent or exfiltrate sensitive information OWASP GenAI Security Project (2025); Liu et al. (2024). For safety-critical deployments, we recommend standard secure-by-design measures (least-privilege tool access, input/output filtering, and monitoring) and adversarial testing. Automated red-teaming benchmarks offer a systematic way to assess safety and refusal robustness Mazeika et al. (2024).
@@
Data contamination and evaluation leakage: Because our training corpus is derived from publicly available instruction-following datasets, there is a non-zero risk of overlap with downstream evaluation items or closely related paraphrases. While we rely on the official SEA-HELM pipeline for standardized comparison, future work should include systematic decontamination checks (e.g., near-duplicate detection between training data and benchmark prompts) to strengthen the validity of reported gains.